OpenStudy (anonymous):

Solve the problem. If csc θ=2, find the value of csc θ+sec(90 degrees-θ).

OpenStudy (anonymous):

\[\csc \theta +\sec (90-\theta)=\csc \theta + \csc \theta=2\csc \theta=2(2)=4\]